OpenStudy (anonymous):

Can someone please check my answer to this equation? |-4b - 8 | + |-1 - b^2 | + 2b^3 ; b = -2 the answer I got is 69, is this correct?

Vocaloid (vocaloid):

nope, not quite, lets break it down step by step

OpenStudy (anonymous):

ok I'm ready

Vocaloid (vocaloid):

|-4b - 8 | = ?

OpenStudy (anonymous):

dont you have to plug -2 into b first?

Vocaloid (vocaloid):

yeah, what do you get when you do that?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

0 sorry

Vocaloid (vocaloid):

actually, yes, 0

OpenStudy (anonymous):

lol i thought so

Vocaloid (vocaloid):

|-4b - 8 | = |8-8| = 0

Vocaloid (vocaloid):

|-1 - b^2 | = ?

OpenStudy (anonymous):

i got -5

Vocaloid (vocaloid):

almost, remember that the absolute value sign means that you remove the negative sign at the end |-1-b^2| = |-1-4| = |-5| = 5, with me so far?

OpenStudy (anonymous):

yes

Vocaloid (vocaloid):

last step, find 2b^3 (be careful with signs!)

OpenStudy (anonymous):

is it -64?

Vocaloid (vocaloid):

yup! now all that's left is 5 - 64 = your answer

OpenStudy (anonymous):

-59

Vocaloid (vocaloid):

good! that's it.

OpenStudy (anonymous):

thank you :)! @Vocaloid